describe the skewness and kurtosis with the distribution graph.
X (income) No. of persons Cumulative frequency (cf)
180 16 16
200 24 40
250 26 66
280 30 96
Answer #1
X No. Of person Cf Xf (X-xbar)^2 f (X-xbar)^3 f (X-xbar)^4 f
180 16 16 2880 48770.3056 -2692608.6 148658919
200 24 40 4800 29753.8584 -1047633.4 36887170.4
250 26 66 6500 5687.3466 84115.8562 1244073.51
280 30 96 8400 60184.323 2695655.83 120738424
Total 96 22580 144395.834 -960470.24 307528588
\small \bar{x}, m_{2},m_{3},m_{4} na na 235.21 1504.12 -10004.898 3203422.79

Skewness,

\small g_{1}=\frac{m_{3}}{m_{2}^{3/2}}=\frac{-10004.898}{1504.12^{3/2}}= -0.172

Kurtosis,

\small g_{2}=\frac{m_{4}}{m_{2}^{2}}=\frac{3203422.79}{1504.12^{2}}= 1.416

Normal 30- Mean 235.2 StDev 38.99 N 96 25 20- Fequency 15 10 - 5 - 0 150 180 210 270 300 240 х

The data is negatively skewed and platykurtic.

****If you have any queries or doubts please comment below. if you're satisfied please give a like. Thank you!